OpenStudy (anonymous):

Find the exact value of the radical expression in simplest form. the square root of 169 − the square root of 50 − the square root of 8

OpenStudy (anonymous):

\[\sqrt{169} - \sqrt{50} - \sqrt{8}\]

OpenStudy (anonymous):



OpenStudy (anonymous):

would it be a 7 or a 3?

OpenStudy (anonymous):

7

OpenStudy (anonymous):